ELECTRONIC GROUND STATE OF 2-CYCLOHEXEN-1-ONE: STRUCTURE AND INVERSION POTEN-TIAL ELUCIDATED BY ULTRAVIOLET CAVITY RINGDOWN SPECTROSCOPY AND DFT CALCULA-TIONS

EMILY J. GILLES, \underline{STEPHEN DRUCKER}, Department of Chemistry, University of Wisconsin-Eau Claire, Eau Claire, WI 54702; JAEBUM CHOO, Department of Chemistry, Hanyang University, Ansan 425-791, Korea; DANIEL AUTREY, Department of Natural Sciences, Fayetteville State University, Fayetteville, NC 28301; MOHAMED RISHARD, JAAN LAANE, Department of Chemistry, Texas A & M University, College Station, TX 77843.

The S_{1}(n, π^{∗}) ← S_{0} cavity ringdown spectrum of 2-cyclohexen-1-one vapor has been recorded in the vicinity of the origin band, which is at 26,089.1 ±0.1cm^{−1}. Observation of hot bands in the spectrum has permitted the determination of several low-frequency fundamentals and overtones in the ground electronic state. The lowest two excited quantum states for the inversion vibration, ν_{39}, were found to be at 99.0 and 197.0 cm^{−1}. Together with previously published far-infrared spectra and vapor-phase Raman spectra, the fundamental frequencies for ν_{37} and ν_{38} have also been established. From observed ν_{39} levels, the barrier to inversion was determined experimentally to be 1900 ±300 cm^{−1}, which is very different from values of 935 and 3379 cm^{−1} previously reported from Raman and far-infrared data, respectively. Density functional calculations carried out in the present work give a barrier value of 2090 cm^{−1} when the B3LYP/6-311 + G(d,p) basis set is used.